The results of a sample of 372 subscribers to Wired magazine shows the time spent using the Internet during the week. Previous surveys have revealed that the population standard deviation is 10.95 hours.

What is the probability that another sample of 372 subscribers spends less than 19.00 hours per week using the Internet?

Descriptive statistics
Column1
count 372
mean 18.54
sample standard deviation 10.81
sample variance 116.75
minimum 2
maximum 38
range 36
standard error of the mean 0.56
confidence interval 95.% lower 17.45
confidence interval 95.% upper 19.64
margin of error 1.10
z 1.96
Answer #1

Here, μ = 18.54, σ = 0.5677 and x = 19. We need to compute P(X <= 19). The corresponding z-value is calculated using Central Limit Theorem

z = (x - μ)/σ
z = (19 - 18.54)/0.5677 = 0.81

Therefore,
P(X <= 19) = P(z <= (19 - 18.54)/0.5677)
= P(z <= 0.81)
= 0.7910